Lucas was 50 years old when his granddaughter was born. Now he is two years more than four times his granddaughter's age. How old is his granddaughter now?
Answer by ankor@dixie-net.com(22740)

Lucas was 50 years old when his granddaughter was born.
Now he is two years more than four times his granddaughter's age.
How old is his granddaughter now?
:
Let L = Lucas's present age
Let d = daughter's present age
From the given information we know:
L = d+50
:
"Now he is two years more than four times his granddaughter's age."
L = 4d + 2
we know L = (d+50), replace L in the above equation
d + 50 = 4d + 2
50 - 2 = 4d - d
48 = 3d
d = 48/3
d = 16 yrs is granddaughter's present age